A projectile's launch speed is 6.70 times its speed at maximum height. Find the launch angle θ0.
1
Expert's answer
2020-02-20T09:53:11-0500

The velocity of the projectile at maximum height


"v=v_0\cos\theta."

Hence


"v_0=6.70v=6.70v_0\cos\theta.""\cos\theta=1/6.70=0.149,"

"\theta=81.4^{\circ}"